Project

General

Profile

Actions

Run Plan 2015/01/15 Controls Status

Following discussion between Chris Rogers, Pierrick Hanlet 10:00 on Tuesday 6th January

Summary

  • Beamline ioc -> finished but problems with integration tests (due to external issues)
  • Beamline state machine -> in progress, due Friday 9th Jan
  • Run control -> not started, no estimate for completion
  • Tracker -> in progress; critical issue is that SVS VME card has been replaced with a CAEN VME card; cannot at present control this even in standalone mode.
  • EMR -> in progress
  • Possible issue with SY527 (TOF HV)

Beamline

  • Beamline ioc
    • Coding work is complete
    • PPS issue prevented operation of dipoles; bobs were not resetting; hence integration test failed
  • Beamline state machine
    • This week maybe
  • Run control
    • Not started don't know
  • Diffuser state machine is near test
  • Diffuser run control is in progress

Fall back on beamline is to run as standalone systems, controlled from the MLCR

Detectors

  • EMR is challenging to complete
    • missing a LV power supply in the hall, one is currently installed
    • HV control is done
    • LV control is in progress
    • VME control is in progress
    • NIM monitoring is in progress
    • Rack switch is in progress
    • temp monitoring - wrong monitors installed, monitoring unlikely to be available in near term
  • Tracker
    • Required new fibre runs, VME issues
    • To run CAEN, need to integrate infrastructure IOC, tracker initialisation, Hanlet monitoring IOC, control IOC
    • Plan to merge control IOC with Hanlet monitoring IOC
    • Significant risk to tracker operation
    • Alan Bross, Paul Rubinov arriving on 15th Jan to 24th Jan

(Shifter training Feb)

  • Chris Heidt will be in charge of the training for the C+M shifter training

Updated by Rogers, Chris almost 9 years ago ยท 2 revisions